A boy standing in the middle of a field, observes a flying bird in the north at an angle of elevation fo 30 degree. and after 2 min, he observes the same bird in the south at an angle of elevation o f60 degree. If the bird flies all along in a straight line at a height of 50root3 m, then its speed in km/h is

Grade:

2 Answers

renu akunuri
11 years ago

ans is 6km/hr

{take tan 60=50root 3/x , tan 30 =50 root 3/y , then add x+y ,this x+y is the distance travelled by it , given time t=2min=2/60 hrs

you will get x+y=200m=0.2 km

velocity =d/t=0.2*60/2 km/hr}

lokesh soni
11 years ago

tan30=50√3/x

1/√3=50√3/x

150=x

√3=50√3/y

y=50

total distance covered =x+y=200m

.2km

time=2/60hr

speed=6km/hr